# Horizon-scale effects in astrophysical black holes from mass-dependent regulators

### v1

@@ Abstract
Typically, regular black holes are described by a regulator $\ell$ that is much smaller than the horizon radius $r_{\text{h}}$; taking the regulator scale large removes the black hole horizon, leaving a compact, horizonless object. This seemingly justifies the well-known lore that horizon scale effects for astrophysical black holes are suppressed by the dimensionless ratio $\ell/r_{\text{h}}\ll 1$. We show explicitly that regular black hole metrics with $\sim 50\%$ corrections near the horizon are, in fact, possible, without prohibiting a range in black hole masses that would conflict with observation. This opens the doors to potentially observable horizon-scale effects for the light deflection and the shadow around such regular, astrophysical black holes.
Introduction.—Astrophysical black holes are emerging as increasingly relevant testing grounds of gravitational physics [1, 2, 3, 4, 5, 6]. Under some assumptions on the energy and pressure of matter, general relativity predicts that such astrophysical black holes—formed during the late stages of a collapsing massive star—necessarily contain singularities [7, 8, 9]. These singularities show up as divergent gravitational tidal forces in the black hole interior, and prevent a complete description of physics in this regime [10]. This is the black hole singularity problem, a major unresolved problem in gravitational physics. Within general relativity, these singularities are evident in the Schwarzschild metric: an exact, spherically symmetric vacuum solution of the Einstein equations, taking the form [11, 12]
@@
Here, $M$ is the black hole mass, $G$ is the gravitational constant, and we set the speed of light to unity ($c=1$). The event horizon is located at $r_{\text{h}}=2GM$, and the singularity is situated at $r=0$, as can be seen from the Kretschmann invariant $R{}_{\mu\nu\rho\sigma}R{}^{\mu\nu\rho\sigma}=48(GM)^{2}/r^{6}$. It is believed that quantum-gravitational effects change the form of the metric in the black hole interior at $r\ll r_{\text{h}}$, eventually leading to an avoidance of the singularity. In the recent years, a variety of regular black hole metrics have been proposed [13, 14, 15, 16, 17]. Typically, they feature a modification of an otherwise singular metric, parametrized by a length scale $\ell>0$, that removes the singularity at $r=0$. These metrics are not solutions of a fundamental gravitational theory, but can help constrain the leading order observational consequences of regular black holes. A characteristic example is the Hayward metric [16],
$F_{\text{H}}(r)$ $=1-\frac{2GM}{r}\frac{r^{3}}{r^{3}+L^{3}}\,,\quad L^{3}=2GM\ell^{2}\,.$ (2)
@@
This metric only describes a black hole if it has an outer horizon located at $F_{\text{H}}(r_{\text{h}})=0$, which requires that $\ell<4GM/(3\sqrt{3})$. The regulator $\ell$ is always smaller than the horizon scale, $\ell<r_{\text{h}}$, if a horizon exists. Conversely, a given regulator scale $\ell$ prevents black holes with a mass below $\sim\ell/G$ [^1] This remains true in dynamical black hole formation [18]. This is referred to as a “mass gap.” Tabletop Eötvös-type experiments can constrain the regulator $\ell$, confirming the Newtonian inverse-square law of gravitation to distances of $\sim 50\,\mu\text{m}$ for test masses of $\mathcal{O}(1\,\text{kg})$. A measurement uncertainty of around $0.5\,\mu\text{m}$ [19] bounds $\ell$ in Eq. (2) to
$\ell\lesssim\mathcal{O}(700\,\text{km})\,.$ (3)
Such large values for $\ell$ are possible since for tabletop masses $2GM/c^{2}\sim 10^{-27}\,\text{m}$, and only $L$ has to be $\lesssim 10\mu\text{m}$. The mass-dependent regulator $L$ removes the mass dependence of the function (2) at small distances $r$, which is known as the limiting curvature condition [20, 21, 22], obeyed by most regular black hole metrics [17]. Since $L$ depends on mass, it will be much larger for astrophysical objects. However, the possibility of observing significant horizon-scale effects in astrophysical regular black holes must overcome two immediate challenges:
  • A)
    Theoretical prejudice: it is usually assumed that $\ell$ is $\mathcal{O}(\ell_{\text{Planck}})$, so that effects at the horizon are minuscule and unobservable.
  • B)
    If $\ell$ is large enough so that horizon-scale effects are maximized for a black hole of mass $M$, then the mass gap prohibits black holes with masses less than $M$, in conflict with observation.
In the absence of a model for the origin of Eq. (2), it is not unreasonable to treat $\ell$ as a free parameter and study the phenomenological consequences. However, challenge B cannot be dispensed with so easily and motivates consideration of modifications of the metric. In this Letter, we propose, for the first time, regular black hole metrics that can feature $\mathcal{O}(50\%)$ corrections at the horizon scale in the mass range of astrophysical black holes, without the problems related to the existence of the mass gap. Unlike the regular black hole metrics described in the literature, the class of regular black holes described here allow the existence of black holes in a regime that is most promising for observational tests of modified gravity. For example, the location of the innermost circular orbit for light (the “photon sphere”), gravitational lensing, and the black hole shadow, are all susceptible to horizon-scale deviations in the astrophysical black hole mass range.
@@
which one may think of as the “black hole condition.” The mass-dependence of this condition leads to strikingly different results compared to the Hayward metric. Namely, non-monotonic functions $\ell^{2}f(\hat{\ell})$ that increase and decrease at intermediate values for $\hat{\ell}$ generate entirely new, astrophysically viable branches for regular black hole metrics that feature both a horizon and a large value of $\ell$. To illustrate this, the reader may verify that the following function,
@@
provides such an example. The black hole condition $\hat{L}^{3}\leq 4/27$ is guaranteed for $GM\gtrsim 1.284\,\ell$, qualitatively similar to the Hayward case. However, black holes also exist for $GM\lesssim 0.195\,\ell$—an entirely new branch. The exciting property of this branch lies in the fact that it constitutes an upper bound on black hole mass. It allows black holes with masses that are less than the regulator scale (in units where $c=G=1$). This provides a means of addressing challenge B, and opens the door to astrophysically large effects without sacrificing the existence of a horizon. The choice of function $f(\hat{\ell})$ defines a new family of regular black holes, as we discuss in some more detail below.
@@
The deviation is monotonically increasing and assumes its maximum value of $50\%$ at $\hat{L}^{3}=4/27$. There, usual regular black hole metrics run into challenge B. The function $f(\hat{\ell})$ in the mass-dependent regulator $\hat{L}$ changes the allowed mass ranges for black holes, so that $\mathcal{O}(50\%)$ horizon-scale effects for astrophysical black holes are allowed; Fig. 1 illustrates the effect of the function $f(\hat{\ell})$. Let us now constrain its properties a bit more:
@@
  • The limiting curvature condition demands $f$ to be a universal constant. If we merely want to prevent trans-Planckian curvatures, we can relax this condition to only hold for large masses $M$ (at fixed $\ell$), and hence $f\sim\hat{\ell}^{p}$ with $p\geq 0$ for ^$\hat{\ell}\rightarrow 0$.
@@
In the black hole domain, as described by Eq. (6), the modification is strongest wherever $\hat{L}^{3}$ reaches its maximum value of $4/27$. For example, in the case where $a=1$, $p=4$ and $b=0$, i.e., Eq. (7), maximal effects are obtained for black hole masses near the boundaries of the allowed regions, $\sim 0.195\ell/G$ and $\sim 1.284\ell/G$, while allowing black holes to exist in both intervals. For $\ell\sim 30\,\text{km}$, these boundaries correspond to $\sim 4$ and $\sim 26$ solar masses, respectively. For realistic astrophysical studies, the forbidden mass interval in between can be modified by a different choice of the function $f(\hat{\ell})$. Note that regular supermassive black holes would not be affected by the modification to the metric. Compared to common regular black hole models, the additional stellar black hole “branches” with substantial horizon-scale effects may be ideal targets for direct astrophysical observation.
@@
and explore observational consequences for astrophysical, solar-mass black holes. For simplicity, we will assume a mass-dependent regulator given by $\{a=1,p=4,b=0\}$.
@@
and the maximum possible deviation from Schwarzschild is again of $\mathcal{O}(5\%)$.
@@
Conclusions.—We have described a new class of regular black hole metrics that can feature potentially observable, astrophysically relevant horizon-scale deviations from the black holes described by general relativity at the percent level. These include, but are probably not limited to, a reduction in the black hole’s apparent size, a smaller photon sphere, and weaker lensing. These effects are made possible across astrophysical black hole mass ranges via a mass-dependent regulator $f(\hat{\ell})$.
Percent-level effects for astrophysical black holes are a novel phenomenon in the field of regular black holes. Usually, large effects are parametrized by a regulator $\ell$ that is thought to be of quantum-gravitational origin, and hence of the order of the Planck length. Effects for astrophysical black holes are thus minuscule, $\ell_{\text{Planck}}/(2GM_{\odot})\sim 10^{-38}$. One may ask: what happens if the fundamental regulator, for some reason, is macroscopic? Then, regular black hole metrics cease to describe black holes with masses below the scale given by the regulator. This is sometimes referred to as a “mass gap,” and leads to unwanted consequences when the regulator is macroscopic.
In this Letter, we resolved this problem by exploring the option of a large regulator $\ell$ that is potentially of the order of kilometers, and constructed a regular black hole metric that still features a horizon for masses at or below $\sim\ell/G$. This is made possible by a mass-dependent regulator
$\hat{L}^{3}=2GM\ell^{2}f\left(\frac{\ell}{2GM}\right)\,,$ (16)
where the function $f$ is a new ingredient. Even in the case of $f=1$ the regulator is mass-dependent [16], mandated by the limiting-curvature condition. It seems like a natural extension to allow for a more general interplay between $L$ and $M$. As we have shown, if this function is non-monotonic in its dimensionless argument, the black hole mass spectrum is enlarged. Previously excluded mass ranges are now accessible and permit black hole geometries featuring horizons, while other ranges remain inaccessible. In a simple case, cf. Eq. (7), we find
$GM\gtrsim 1.284\ell\,,\quad GM\lesssim 0.195\,\ell\,.$ (17)
While the first constraint constitutes a mass gap, the second condition is qualitatively new. This, to the best of our knowledge, is the first example of a “black hole band spectrum” in the context of regular black holes.
One may wonder about the dynamical aspects of these black holes and their non-trivial band structures. For example, is it possible to have two black holes from the low-mass part of the spectrum, make them collide, and end up in the regime where no black holes are allowed? And, in that case, what is the nature of the horizonless end product after a collision of these black holes? It is possible that instabilities of this compact, horizonless object lead to emission of gravitational waves that only terminates once the object is again inside the lower mass range and a horizon has formed. In the presence of more involved regulators $f(\hat{\ell})$ one could have multiple, disconnected bands for these types of black holes, giving rise to a multitude of transitions between bands, accompanied by a characteristic amount of gravitational wave emission. While certainly interesting, we will leave dynamical aspects to future work.
@@
We also briefly comment on the issue of supermassive black holes. Since their mass scale exceeds that of astrophysical black holes up to a power of $10^{9}$ [23], the effects described in this Letter are suppressed by roughly this factor. This is desirable, since the shadows of supermassive black holes are directly observable [4], and large deviations from general relativity are likely ruled out at that scale.
Last, one may wonder if these deviations, if detected, constitute modified, classical gravity, or contain traces of quantum aspects of gravity. The appearance of $GM$ in the regulator $L$, even in the case of the unmodified Hayward metric, is necessitated by the limiting curvature condition. This condition states that irrespective of its mass, the maximally allowed curvature of an object cannot exceed the Planckian curvature $\mathcal{R}\sim 1/\ell_{\text{Planck}}^{2}$. In other words, the maximum curvature needs to be independent of the object’s mass. This condition (sometimes also called conjecture) is rooted in quantum-gravitational considerations. Therefore, regulators with an unconventional mass dependence, like the ones described in this Letter, may very well one day traced back to a theory of quantum gravity.
Acknowledgements.—We are grateful for support by the National Science Foundation under grant no. PHY-2112460.

### v3

# Note on black holes with kilometer-scale ultraviolet regulators

@@ Abstract
Regular black hole metrics involve a universal, mass-independent regulator that can be up to $\mathcal{O}(700\,\text{km})$ while remaining consistent with terrestrial tests of Newtonian gravity and astrophysical tests of general relativistic orbits. However, for such large values of the regulator scale, the metric describes a compact, astrophysical-mass object with no horizon rather than a black hole. We note that allowing the regulator to have a nontrivial mass dependence preserves the horizon, while allowing large, percent-level effects in black hole observables. By considering the deflection angle of light and the black hole shadow, we demonstrate this possibility explicitly.
Introduction.—Astrophysical black holes are emerging as increasingly relevant testing grounds of gravitational physics Will:2014kxa ; Berti:2015itd ; Barack:2018yly ; EventHorizonTelescope:2019dse ; LIGOScientific:2020tif ; Horowitz:2023xyl . Under some assumptions on the energy and pressure of matter, general relativity predicts that black holes formed during the late stages of a collapsing massive star necessarily contain singularities Penrose:1964wq ; Hawking:1970zqf ; Hawking:1973uf . These singularities show up as divergent gravitational tidal forces in the black hole interior, and prevent a complete description of physics in this regime Geroch:1968ut . This is the black hole singularity problem, a major unresolved problem in gravitational physics. Within general relativity, these singularities are evident in the Schwarzschild metric: an exact, spherically symmetric vacuum solution of the Einstein equations, taking the form Schwarzschild:1916 ; Droste:1917
@@
Here, $M$ is the black hole mass, $G$ is the gravitational constant, and we set the speed of light to unity ($c=1$). The event horizon is located at $r_{\text{h}}=2GM$, and the singularity is situated at $r=0$, as can be seen from the Kretschmann invariant $R{}_{\mu\nu\rho\sigma}R{}^{\mu\nu\rho\sigma}=48(GM)^{2}/r^{6}$. It is believed that quantum-gravitational effects change the form of the metric in the black hole interior at $r\ll r_{\text{h}}$, eventually leading to an avoidance of the singularity. In the recent years, a variety of regular black hole models have been proposed Bardeen:1968 ; Dymnikova:1992ux ; Bonanno:2000ep ; Hayward:2005gi ; Simpson:2018tsi ; Frolov:2021vbg ; see also the review Frolov:2016pav . Typically, they feature a modification of an otherwise singular metric, parametrized by a length scale $\ell>0$, that removes the singularity at $r=0$. These metrics are not solutions of a fundamental gravitational theory, but can help constrain the leading order observational consequences of regular black holes. A characteristic and popular model is the Hayward metric Hayward:2005gi ,
$F(r)$ $=1-\frac{2GM}{r}\frac{r^{3}}{r^{3}+L^{3}}\,,\quad L^{3}=2GM\ell^{2}\,.$ (2)
@@
For non-zero values of the parameter $L$, the metric in Eq. (2) describes a smooth and finite gravitational field around $r=0$, while for $L=0$ one recovers the Schwarzschild solution. The parameter $L$ is the product of a universal length scale $\ell$ and a factor of $GM$ that removes the mass dependence of the function (2) at small distances $r$. This universal behavior at small distances is known as the limiting curvature condition and is motivated by various high-energy properties of gravity Markov:1982 ; Markov:1984 ; Polchinski:1989 ; Mukhanov:1991zn . Hence we shall refer to $\ell$ (rather than $L$) as the fundamental, short-distance or “ultraviolet” (UV) regulator.
Let us now constrain both $L$ and $\ell$. Eötvös-type experiments confirm the Newtonian inverse-square law of gravitation to distances of $\sim 50\,\mu\text{m}$ for test masses of $\mathcal{O}(1\,\text{kg})$. A measurement uncertainty of around $0.5\,\mu\text{m}$ Lee:2020zjt bounds $L$ and $\ell$ by
$L\lesssim\mathcal{O}(10\,\mu\text{m})\,,\qquad\ell\lesssim\mathcal{O}(700\,\text{km})\,.$ (3)
[…]
In the absence of a model for the origin of Eq. (2), it is well motivated to treat $\ell$ as a free parameter and study the phenomenological consequences. The surprisingly weak bound in (3) implies that $\ell$ can be comparable to the horizon size of astrophysical black holes. When this is the case, regular black hole metrics can differ significantly in the near-horizon region from their Schwarzschild counterparts. The possibility of striking observational consequences, however, cannot be realized: the metric (2) ceases to describe a black hole if $GM<(3\sqrt{3}/4)\ell$, preventing black holes with masses below $\sim\ell/G$ [^1] This remains true in dynamical black hole formation Frolov:2015bta . Importantly, this is not a fluke of the Hayward metric, but a common feature of all known regular black hole metrics, usually referred to as a “mass gap.” See Table 1 for a summary of mass gaps for several well-known regular black hole models. To the best of our knowledge, no known regular black hole model can simultaneously accommodate a large regulator $\ell$ and allow the existence of a black hole horizon across a realistic, astrophysical mass range.
In this Note, we show how one may obtain a different outcome by allowing the regulator $\ell$ to have a nontrivial mass dependence. Put in another way, the quantity $L^{3}$ in Eq. (2) is conventionally assumed to be mass dependent, but the present work explores the phenomenological implications of relaxing the assumption that this dependence is linear. In doing so, we obtain regular black hole metrics without the problems related to mass gaps, featuring $\mathcal{O}(50\%)$ corrections at the horizon scale in mass ranges associated with astrophysical black holes. Given our assumptions about the mass dependence of the regulator, the motivation for considering kilometer-scale values of $\ell$ is similar to the motivation for considering large compactification radii in extra-dimensional scenarios. When the large extra dimensions idea was first proposed Dienes:1998vg ; Arkani-Hamed:1998jmv , there was no theoretical derivation that indicated a priori that one or more compactification radii should be at the millimeter or the inverse-TeV scale. The proposal was made to challenge a theoretical prejudice, namely that all compactification scales should be set by the Planck scale. Dispensing with this preconception led to significant phenomenological benefits, most notably the elimination of the hierarchy problem. In the present context, there is no theoretical derivation that indicates a priori that the gravitational regulator $\ell$ should be kilometer-scale in size. Our scenario challenges a theoretical prejudice that the regulator $\ell$ should be set by the Planck scale. Dispensing with this preconception also leads to significant phenomenological benefits, most notably the amelioration of the black hole mass gap problem and the possibility of observable horizon-scale effects. The location of the innermost circular orbit for light (the “photon sphere”), gravitational lensing, and the black hole shadow, are all susceptible to horizon-scale deviations in our scenario. The purpose of this Note is to illustrate these points, which have not been made previously in the literature. Nevertheless, fully realistic black holes must also address other issues, such as geodesic completeness and stability, and take into account black hole angular momentum. We discuss these briefly in the final section, as motivation for future work.
Regular black hole model Mass gap Ref.
Bardeen $GM\geq 1.30\ell$ Bardeen:1968
Dymnikova $GM\geq 0.88\ell$ Dymnikova:1992ux
Bonanno–Reuter ($\gamma=9/2$) $GM\geq 3.50\ell$ Bonanno:2000ep
Hayward $GM\geq 1.30\ell$ Hayward:2005gi
Simpson–Visser $GM\geq 0.50\ell$ Simpson:2018tsi
Frolov ($n=1$) $GM\geq 0.98\ell$ Frolov:2021vbg
Table 1: Mass gaps for several regular black hole models.
@@
which one may think of as the “black hole condition.” As in Eq. (4), and henceforth, a hat indicates that the given quantity has been made dimensionless by dividing by an appropriate power of $2GM$. The mass-dependence of this condition leads to strikingly different results compared to the Hayward metric. Namely, non-monotonic functions ^$\hat{\ell}^{2}f(\hat{\ell})$ that increase and decrease at intermediate values for $\hat{\ell}$ generate entirely new, astrophysically viable branches for regular black hole metrics that feature both a horizon and a large value of $\ell$. To illustrate this, the reader may verify that the following function,
@@
provides such an example. The black hole condition $\hat{L}^{3}\leq 4/27$ is guaranteed for $GM\gtrsim 1.28\,\ell$, qualitatively similar to the Hayward case. However, black holes also exist for $GM\lesssim 0.20\,\ell$—an entirely new branch. It is significant that this branch constitutes an upper bound on black hole mass, one that allows black holes with masses that are less than the regulator scale (in units where $c=G=1$). This solves the problem of a disappearing horizon for large regulator scales, so that large effects on black hole observables can be consistently obtained. The choice of function $f(\hat{\ell})$ defines a new family of regular black holes, as we discuss in some more detail below.
@@
The deviation is monotonically increasing and assumes its maximum value of $50\%$ at $\hat{L}^{3}=4/27$. There, usual regular black hole metrics cease to describe black holes and instead become horizonless. The function $f(\hat{\ell})$ in the mass-dependent regulator $\hat{L}$ changes the allowed mass ranges for black holes, so that $\mathcal{O}(50\%)$ horizon-scale effects for astrophysical black holes are allowed; Fig. 1 illustrates the effect of the function $f(\hat{\ell})$. Let us now constrain its properties a bit more:
@@
  • Close to the origin the metric and its Kretschmann scalar behave as
    $\begin{split}F(r\rightarrow 0)&=1-\frac{1}{f}\frac{r^{2}}{\ell^{2}}+\mathcal{O}(r^{5})\,,\\ K(r\rightarrow 0)&=\frac{1}{f^{2}}\frac{24}{\ell^{4}}+\mathcal{O}(r^{3})\,,\end{split}$ (10)
    where $K\equiv R{}_{\mu\nu\rho\sigma}R{}^{\mu\nu\rho\sigma}$ and we suppressed the argument of $f=f(\hat{\ell})$ for brevity. The limiting curvature condition demands $f$ to be a universal constant to avoid any mass dependence in the maximum curvature. However, if we are willing to allow for a mild mass dependence that still prevents trans-Planckian curvatures, we can relax this condition and merely demand that $f$ approaches a constant as $M$ is taken large with $\ell$ fixed, or equivalently $f(\hat{\ell})\rightarrow\mbox{constant}$ as ^$\hat{\ell}\rightarrow 0$.
@@
In the black hole domain, as described by Eq. (6), the modification is strongest wherever $\hat{L}^{3}$ reaches its maximum value of $4/27$. For example, in the case where $a=1$, $p=4$ and $b=0$, i.e., Eq. (7), maximal effects are obtained for black hole masses near the boundaries of the allowed regions, $\sim 0.20\ell/G$ and $\sim 1.28\ell/G$, while allowing black holes to exist in both intervals. For $\ell\sim 30\,\text{km}$, these boundaries correspond to $\sim 4M_{\odot}$ and $\sim 26M_{\odot}$, respectively. The forbidden mass interval in between can be modified by a different choice of the function $f(\hat{\ell})$. Note that regular supermassive black holes are not affected by the modification to the metric. Should our qualitative results persist when angular momentum is introduced, the additional stellar black hole “branches” with substantial horizon-scale effects suggest new and potentially interesting targets for direct astrophysical observation.
@@
and explore observational consequences for solar-mass black holes. For simplicity, we will assume a mass-dependent regulator given by $\{a=1,p=4,b=0\}$.
@@
and the maximum possible deviation from Schwarzschild is again of $\mathcal{O}(5\%)$. In the future, it would be interesting to extend these studies similar to Ref. Carballo-Rubio:2018jzw .
Conclusions.—We have described a new class of regular black hole metrics that do not imply a lower bound on the black hole mass below which one would obtain a compact horizonless object. This is made possible by replacing the regulator $\ell$ by a mass-dependent function, entering the metric Eq. (2) via
$L^{3}=2GM\ell^{2}f\left(\frac{\ell}{2GM}\right)\,,$ (17)
@@
where the function $f$ is the new ingredient. Previously excluded mass ranges are now accessible and permit black hole geometries featuring horizons, while other limited ranges remain inaccessible, without necessarily implying conflict with observation. The mass dependence allows the regulator $\ell$ and the horizon scale to be comparable, so that horizon-scale deviations from the black holes described by general relativity at the percent level can now be obtained. These include, but are probably not limited to, a reduction in the black hole’s apparent size, a smaller photon sphere, and weaker lensing.
@@
Effects on supermassive black holes are negligible given an appropriate choice of $f$. Since their mass scale exceeds that of astrophysical black holes up to a power of $10^{9}$ Bambi:2017iyh , the effects described in this Letter are suppressed by roughly this factor. This is desirable, since the shadows of supermassive black holes are directly observable EventHorizonTelescope:2019dse , and large deviations from general relativity are likely ruled out at that scale. In fact, the observations presently cannot distinguish between black holes from Einstein gravity and modified, regular black hole metrics EventHorizonTelescope:2022xqj ; Vagnozzi:2022moj .
The purpose of this Note has been to illustrate the potential usefulness of mass-dependent black hole regulators in the simplest setting, through modification of the form of well-known Hayward metric. However, a fully realistic application will have to address a number of issues that we reserve for future work:
$\bullet$ Geodesic Completeness and Stability. In the recent years, it has been pointed out that several regular black hole models are geodesically incomplete Bambi:2016wdn ; Carballo-Rubio:2019fnb ; Zhou:2022yio or may suffer from instability issues Carballo-Rubio:2018pmi related to mass inflation Poisson:1989zz at their inner horizons. While our modification does not remedy this behavior, we emphasize that our proposed mass-dependent regulator can be applied to any regular black hole model, including models that have improved behavior Carballo-Rubio:2022kad .
$\bullet$ Angular Momentum. Realistic astrophysical black holes have angular momentum. The static metric presented in this Note can be endowed with angular momentum following the well-established procedure presented in Ref. Azreg-Ainou:2014pra ; alternatively, one may replace the regulator in known rotating regular black hole metrics Simpson:2021dyo directly with the mass-dependent regulator proposed in this Note.
$\bullet$ Energy Conditions. The absence of singularities in regular black hole models is typically accompanied with the violation of energy conditions, which has been studied in great detail elsewhere Zaslavskii:2010qz ; Balart:2014jia . In the context of the proposed model—where the radial dependence of the metric coincides with that of the Hayward metric—the only difference emerges through the re-scaled regulator:
$L=(2GM)^{1/3}\ell^{2/3}f^{1/3}\,.$ (18)
Recall that for the astrophysical mass range, kilometer-scale fundamental regulators $\ell$ would give an $L$ that is also of the order of kilometers—were it not for the modification function $f$ that can be chosen to reduce the total regulator $L$. Since the scale $L$ sets the violation length scale for the energy conditions, in the proposed model the violation will be constrained to that modified range, which, for astrophysically relevant black holes, is smaller than in the pure Hayward case. In this case, this model therefore shrinks the effective regions of energy condition violations. However, we also note that energy conditions, in their standard form, only apply to general relativity, wherein the energy-momentum tensor is algebraically related to the Ricci curvature. In modified gravity theories (or, for example, an effective theory of quantum gravity that may dynamically generate regular black hole metrics like the one discussed in this Note) this algebraic relation may no longer be true. In the absence of an action principle and equations of motion, such questions cannot be unambiguously answered. [^2] This also applies to stability issues related to black hole thermodynamics, since the latter as a semiclassical approach can only be defined in the presence of an action.
$\bullet$ Gravitational Waves. Since the discussion in this Letter is limited to time-independent metrics, one may wonder about the dynamical aspects of these black holes and their non-trivial band structures (i.e., allowed mass intervals). For example, is it possible to have two black holes from the low-mass part of the spectrum, make them collide, and end up in the regime where no black holes are allowed? And, in that case, what is the nature of the horizonless end product after a collision of these black holes? It is well known that such compact, horizonless object can feature instabilities Cunha:2022gde . This may then lead to emission of gravitational waves that only terminates once the object is again inside the lower mass range and a horizon has formed. In the presence of more involved regulators $f(\hat{\ell})$ one could have multiple, disconnected bands for these types of black holes, giving rise to a multitude of transitions between bands, accompanied by a characteristic amount of gravitational wave emission. While certainly interesting, we will leave dynamical aspects to future work.
Lastly, we point out that both the unmodified Hayward metric and the metrics proposed in this Note satisfy the limiting curvature condition Markov:1982 ; Markov:1984 ; Polchinski:1989 , which is rooted in quantum-gravitational considerations. Thus, one might hope that regulators with an unconventional mass dependence may one day be traced back to an underlying theory of quantum gravity. [^3] One might view the unconventional mass dependence of the regulator in Ref. Bonanno:2000ep as a step in this direction.
Acknowledgements.—We are grateful for support by the National Science Foundation under grant no. PHY-2112460. J.B. acknowledges support as a Fellow of the Young Investigator Group Preparation Program administered by the state of Baden–Württemberg (Germany) and the Karlsruhe Institute of Technology.